Martin, Inc., is considering the development of a subsidiary in Sydney, Australia that would manufacture and...

  • Martin, Inc., is considering the development of a subsidiary in Sydney, Australia that would manufacture and sell high quality guitars on the local Australian Market.
  • As one of Martin’s financial managers, you have asked the finance, marketing and manufacturing departments to provide you with all of the relevant input so you can perform a capital budgeting analysis to determine whether to undertake this project or not.
  • In addition, a contingent of Martin executives have met with Australian government officials in Sydney to discuss the proposed subsidiary.
  • The project would end in 5 years. All of the relevant information that you need to perform this capital budgeting analysis follows here.
  • Initial investment: A$12,500,000 million (A$ = Australian dollars)
  • Price and consumer demand:

Year 1 and 2: 20,000 units @ A$650/unit

Year 3: 30,000 units @ A$760/unit

Year 4: 40,000 units @ A$850/unit

Year 5: 40,000 units @ A$880/unit

  • Costs

Variable costs: Years 1 & 2 A$370/unit, Year 3 A$445/unit, Years 4 and 5 A$600/unit

Fixed costs: A$2,000,000 per year ( A$1,000,000 lease and A$1,000,000 other fixed expense)

  • Depreciation A$1,000,000 per year
  • Tax laws: 30% income tax
  • Remitted funds: 15% withholding tax on remitted funds
  • Exchange rates: Spot exchange rate of $US 0.65 for Australian dollar
  • Salvage values: A$6 million
  • Required rate of return: 18%

Given all of the preceding information, please create an Excel spread sheet that shows the year by year amounts that you have taken into consideration in your capital budget analysis and decide whether or not you will undertake this project.

Show your work; indicate your decision (do we do the project or not); and briefly explain
Answer #1
Present Value(PV) of Cash Flow:
(Cash Flow)/((1+i)^N)
i=discount rate=Required Return=18%=0.18
N=Year   of Cash Flow
Accumulated depreciation at end of 5 years(A$)               5,000,000 (5*1 million)
Book Value at end of 5 years               7,500,000 (12500000-5000000)
Salvage Value (A$)               6,000,000
Loss on Salvage               1,500,000
Tax Saving on Loss=30%*1500000                   450,000
After Tax Salvage Cash Flow (A$)               6,450,000 (6000000+450000)
N Year 0 1 2 3 4 5
A Initial investment(A$)          (12,500,000)
b Demand in Units                  20,000                  20,000                  30,000                    40,000                40,000
c Sales Price per unit(A$)                        650                        650                        760                          850                      880
d=b*c Annual Revenue(A$)          13,000,000          13,000,000          22,800,000            34,000,000        35,200,000
e Variable Cost per unit (A$)                        370                        370                        445                          600                      600
f=b*e Total Variable Costs(A$)            7,400,000            7,400,000          13,350,000            24,000,000        24,000,000
g=d-f Contribution Margin(A$)            5,600,000            5,600,000            9,450,000            10,000,000        11,200,000
h Fixed Costs(A$)            2,000,000            2,000,000            2,000,000              2,000,000          2,000,000
i Depreciation expenses(A$)            1,000,000            1,000,000            1,000,000              1,000,000          1,000,000
j=g-h-i Before tax Operating Income(A$)            2,600,000            2,600,000            6,450,000              7,000,000          8,200,000
k=j*(1-0.3) After Tax Operating Income (A$)            1,820,000            1,820,000            4,515,000              4,900,000          5,740,000
l Add:Depreciation (Non Cash expense)            1,000,000            1,000,000            1,000,000              1,000,000          1,000,000
X=k+l Annual Operating Cash Flow(A$)            2,820,000            2,820,000            5,515,000              5,900,000          6,740,000
Y Terminal Salvage Cash Flow(A$)          6,450,000
CF=A+X+Y Net Cash Flow          (12,500,000)            2,820,000            2,820,000            5,515,000              5,900,000        13,190,000
p Exchange Rate (USD per AUD) 0.65                       0.65                       0.65                       0.65                         0.65                     0.65
q=CF*p Cash Flow in USD before withholding tax -8125000 1833000 1833000 3584750 3835000 8573500
r Withholding tax 0.15 0.15 0.15 0.15 0.15
s=q*(1-r) Cash Flow in USD after withholding tax -8125000 1558050 1558050 3047038 3259750 7287475 SUM
PV=s/(1.18^N) Present Valure             (8,125,000)            1,320,381            1,118,967            1,854,521              1,681,343          3,185,422          1,035,635 USD
NPV=Sum of PV Net Present Value $1,035,635 USD
Yes, we should undertake this project
NPV is positive